# LLM Output Homogenization is Task Dependent

### v1

@@ Abstract
A large language model can be less helpful if it exhibits output response homogenization. But whether two responses are considered homogeneous, and whether such homogenization is problematic, both depend on the task category. For instance, in objective math tasks, we often expect no variation in the final answer but anticipate variation in the problem-solving strategy. Whereas, for creative writing tasks, we may expect variation in key narrative components (e.g. plot, genre, setting, etc), beyond the vocabulary or embedding diversity produced by temperature-sampling. Previous work addressing output homogenization often fails to conceptualize diversity in a task-dependent way. We address this gap in the literature directly by making the following contributions. (1) We present a task taxonomy comprised of eight task categories that each have distinct conceptualizations of output homogenization. (2) We introduce task-anchored functional diversity to better evaluate output homogenization. (3) We propose a task-anchored sampling technique that increases functional diversity for task categories where homogenization is undesired, while preserving homogenization where it is desired. (4) We challenge the perceived existence of a diversity-quality trade-off by increasing functional diversity while maintaining response quality. Overall, we demonstrate how task dependence improves the evaluation and mitigation of output homogenization.
@@ 1 Introduction
Existing approaches to reducing output homogenization rarely take task dependence into account. Several recent works propose methods that promote diversity in the alignment process or when sampling outputs at inference-time. However, these studies often fail to conceptualize diversity in a task-specific way. For example, some methods aim to increase token-level entropy or embedding-space variation in alignment (Chung et al., 2025; Lanchantin et al., 2025a; Slocum et al., 2025; Li et al., 2025), while others promote diversity of viewpoints and perspectives when sampling multiple outputs (Wang et al., 2025b; Zhang et al., 2025a, b). Without a task-dependent approach, such methods may (1) fail to encourage diversity that is meaningful for a task, and/or (2) undesirably reduce homogenization in tasks where it is desired. We address this gap in the literature directly.
We introduce a task-anchored framework to evaluate and mitigate output homogenization. We build on the notion of functional diversity (Zhang et al., 2025b; Shypula et al., 2025), which asks whether a user would perceive two responses as meaningfully different for a given task. We argue that LLMs should be able to conceptualize functional diversity based on the task category. Consider the stakes: if a model wrongly conceptualizes functional diversity for a task that mimics an encyclopedia inquiry, the model could misrepresent historical events in an attempt to naively reduce homogenization. Conversely, if a model wrongly conceptualizes functional diversity for a creative writing task, the model might repeat the same story arc no matter how many times a user asks the model to tell a story. We argue that task dependence should be incorporated into the way we address homogenization. Our contributions are as follows.
@@
    We introduce task-anchored the functional diversity to better evaluate output homogenization (§ 3.2). In our experiments, we compare to more general diversity metrics which are not task dependent (vocabulary and embedding differences). The results show that these general metrics fail to capture task-dependent diversity. Our task-anchored metric offers an alternative evaluation approach for future studies of output homogenization.
@@
    We propose a task-anchored sampling technique to increase functional diversity (§ 3.3), improving on previous sampling methods to promote diversity (Zhang et al., 2025a, b). Figure 1 offers a high-level illustration of our approach. We leverage our taxonomy to instruct models with task-dependent notions of diversity. Our approach increases functional diversity for task categories where homogenization is undesired, while preserving homogenization where it is desired (Figure 2).
@@
    We challenge the perceived existence of a diversity-quality trade-off (a common narrative in the literature) by adopting a quality measure (Lin et al., 2025; Wei et al., 2025) that also accounts for task-specific variations in quality. Figure 3 shows that the diversity-quality tradeoff prevalent in previous studies may simply be the result of mis-conceptualizing both diversity and quality. Our evaluation framework corrects both.
@@ 2.1 Homogenization in Aligned Models
Task Dependence    Several works show that aligned LLMs exhibit output homogenization across a variety of tasks, such as creative writing (Moon, 2024; Wu et al., 2025), political discussions (Durmus et al., 2023; Santurkar et al., 2023), and math problem-solving (Slocum et al., 2025). Zhang et al. (2024, 2025b) further show that in question-answering, models often produce the same answer, even when the question is underspecified and multiple valid answers exist. These studies often evaluate homogenization in specific task domains, suggesting that problematic notions of homogenization are task-dependent. Our proposed taxonomy compares a variety of these task categories.
@@
Representation Concerns    In certain tasks, homogeneous outputs may raise concerns about representation. The literature on pluralistic alignment (Sorensen et al., 2024; Chen et al., 2024; Zhang et al., 2025a) highlights representational harms, particularly when users seek advice or opinions from LLMs. However, pluralistic alignment discussions tend to operate in contexts where representation or diversity is presumed to be desirable. These discussions should recognize the task dependent nature of pluralism or representation, as we discuss in this work.
@@
Causes    There are many causes of output homogenization, such as limited diversity in training data and model design choices (Zhang et al., 2025a; Fazelpour and Fleisher, 2025). In particular, the alignment process is well-known to amplify homogenization in LLM outputs (Kirk et al., 2024; Lanchantin et al., 2025a). Models are typically aligned using methods such as Reinforcement Learning with Human Feedback (RLHF) (Ziegler et al., 2019) or Direct Preference Optimization (DPO) (Rafailov et al., 2023). These methods involve training on a dataset of pairwise preferences $\{(x,y^{+},y^{-})\}$, where $x$ is a prompt, $y^{+}$ is a preferred response, and $y^{-}$ is a dispreferred response. When there are conflicting preferences in the training data, such that both $y^{+}\succ y^{-}\,|\,x$ and $y^{-}\succ y^{+}\,|\,x$ coexist, the RLHF and DPO objectives implicitly reward putting all sequence-level probability on the majority preference (Slocum et al., 2025; Yao et al., 2025). Preference pairs with larger semantic differences also exert a stronger influence on the behavior of the aligned model (Chung et al., 2025; Shen et al., 2024). While this line of research is important, the present work focuses on how output homogenization should be conceptualized, not why it occurs.
@@
Outcome Homogenization   Another type of homogenization occurs when multiple models produce similar outputs (Kim et al., 2025; Wenger and Kenett, 2025). For example, outcome homogenization in decision-making refers to when individuals receive similar decisions from separate AI models (Bommasani et al., 2022; Jain et al., 2024b). In this work, we focus on the single-model case and do not deal directly with homogenization across different models. But reducing homogenization within a single model likely affects homogenization across models (Jain et al., 2024a).
@@ 2.2 Diversity-Promoting Methods
Alignment Methods   A growing body of literature explores methods to reduce homogenization in aligned LLMs. Several studies propose modifying the alignment process, either by altering the construction of preference datasets or by adjusting alignment objectives (Lanchantin et al., 2025a; Slocum et al., 2025; Chung et al., 2025). All of these methods substantially increase diversity during alignment, as measured by token-level entropy or embedding-space variation. However, we highlight how these metrics may not capture meaningful, task-dependent notions of diversity.
Inference-Time Methods   While most evaluations of homogenization examine temperature-sampled outputs, a few studies explore prompt-based strategies to explicitly sample diverse outputs at inference-time. For example, Zhang et al. (2025a) use a system prompt that explicitly tells LLMs to generate $k$ responses in a single output that represent “diverse values.” Zhang et al. (2025b) propose in-context regeneration, where models are prompted to produce a different response while retaining all previous responses in the conversation context. Other works also use implicit techniques, such as persona-based or multilingual prompting (Wang et al., 2025a, b). Our work improves inference-time methods by explicitly clarifying the notion of “diversity” in model instructions.
@@ 3 Framework
Task Category Task Definition Example Task Functional Diversity Reward Type
A Well-Specified
@@
Tasks with a single verifiable
correct answer
What is the largest Spanish-speaking country? None Verifiable
B Underspecified
@@
Tasks with multiple verifiable
correct answers
Name one Spanish-speaking country. Different correct
@@
Verifiable (Multiple)
C Random
@@
Tasks that involve randomizing over a set of options Roll a make-believe 6-sided dice. Different pseudo-
@@
Verifiable (Multiple)
D Problem-Solving
@@
Tasks to solve a problem with a
verifiable solution
How many positive whole-number divisors does 196 have? Different solution
strategies
@@
E Problem Solving or
Design Subjective
Tasks to solve a problem with many verifiable solutions Design a room that minimizes energy consumption while
maintaining comfort.
Different solutions Partially Verifiable
F Encyclopedia
@@
Tasks to provide information about real world societies, traditions, events where there are credible references Why is Isaac Newton famous? Different factual
@@
Partially Verifiable
G Creative Writing Tasks that require creative
@@
Non-Verifiable
H Advice or Opinions Tasks that solicit advice, opinions or feedback on specific topics/scenarios What is a good Mother’s day gift? Different views or perspectives Non-Verifiable
@@ 3.1 Task Taxonomy
We begin by outlining the task categories used in our task-anchored framework (Table 1). Each of the $8$ categories are distinguished by their conceptualization of output homogenization. The first four categories (A, B, C, D) capture prompts that elicit verifiable solution(s) that might be considered objective in nature, yet may still have more than one verifiable answer or explanation [^1] We view objective/subjective and underspecified/well-specified not as discrete categories, but as spectra that task categories span. While the terms “objective” and “subjective” invite philosophical debate, such discussions are beyond this paper’s scope. Likewise, we use “underspecified” and “well-specified” loosely, as clarifying terms, not precise definitions of the answer spaces.. The second four categories (E, F, G, H) capture prompts that elicit more open-ended solution(s) that may be considered non-verifiable or only have partially verifiable components. Our taxonomy offers a more granular categorization of reward verifiability than the binary distinction (verifiable vs non-verifiable) present in the literature (Lambert et al., 2024; Lanchantin et al., 2025b). Our task categories capture many real-world LLM use cases identified in recent work (Tamkin et al., 2024; Chatterji et al., 2025). Though our taxonomy is non-exhaustive, we illustrate how task categories are a useful mechanism to appropriately conceptualize output homogenization.
@@
Each task category further corresponds to a specific type of response variation or functional diversity. Previous works define two responses to be functionally diverse if a user would perceive them to be meaningfully different (Zhang et al., 2025b; Shypula et al., 2025). Our task categories clarify: in what ways could responses be meaningfully different for a given task? For example, in Problem Solving Objective (category D) tasks, functional diversity is in solution strategies, not in the final answer. Whereas, in Creative Writing (category G), functional diversity is in the key creative elements (e.g. plot, genre, setting, etc.), not just in character names or vocabulary choices.
@@ 3.2 Evaluating Task-Anchored Diversity
Next, we formalize task-anchored functional diversity. We let $\mathcal{P}$ denote the set of possible input prompts or tasks and we let $\mathcal{Y}$ denote the set of possible outputs (e.g., sequences of tokens). We adopt the simple notation that a language model $\mathcal{M}$ is a stochastic function $\mathcal{M}:\mathcal{P}\to\mathcal{Y}$ that maps each prompt $p\in\mathcal{P}$ to an output $y\in\mathcal{Y}$. Each prompt $p$ is associated with a task category $c(p)\in\mathcal{T}$ as defined in Table 1. For a given prompt $p$, we assume $d(p,y_{a},y_{b})\in[0,1]$ to be a metric indicates whether two responses $y_{a}$ and $y_{b}$ differ. To specify functional diversity, we anchor the definition of $d(p,y_{a},y_{b})$ on the task category $c(p)\in\mathcal{T}$.
Definition 3.1 (Task Anchored Functional Diversity).
@@
For example, consider whether two responses are functionally diverse in the Problem-Solving Objective task (category D). Here, $d(p,y_{a},y_{b})$ represents whether responses have different solution strategies, and assumes the single verifiable answer to be the same. In practice, evaluating functional diversity requires human annotation or LLM-judges.
Response diversity can also be evaluated with general diversity metrics. By general, we mean that the metric does not reference or depend on a predefined task category. Previous studies adopt two common metrics, which we define below. Vocabulary diversity quantifies the extent to which two responses use different vocabulary where higher values mean more unique words or less words shared. Embedding diversity measures the difference in semantic content according to cosine distance in an embedding vector space where higher values mean more semantic difference.
Definition 3.2 (Vocabulary Diversity).
Definition 3.3 (Embedding Diversity).
$d_{\mathrm{embed}}(y_{a},y_{b}):=1-\cos(e(y_{a}),e(y_{b})),$
3.3 Promoting Task-Anchored Diversity
We introduce a task-anchored sampling technique which modifies existing prompt-based methods for promoting diversity (c.f. Figure 1). Prompt-based sampling strategies are inference-time methods to generate multiple responses. We focus on two existing methods: system prompt sampling, which generates multiple responses in a single generation (Zhang et al., 2025a), and in-context regeneration, which iteratively generates multiple responses (Zhang et al., 2025b). Both these methods instruct the model to generate “different” or “diverse” responses. We modify these methods by clarifying in the instruction what is meant by “different” or “diverse”, using the functional diversity concepts in our task taxonomy (Table 2). To reduce homogenization at inference-time, the model could sample over these responses, or choose a response based on other alignment criteria.
@@ 3.3 Promoting Task-Anchored Diversity
In-Context
Regeneration
Can you generate a
@@
Table 2: Prompt-Based Sampling Strategies. We modify prompt-based sampling methods in previous works (Zhang et al., 2025a, b) to promote task-anchored functional diversity.

### v2

@@ Abstract
A large language model can be less helpful if it exhibits output response homogenization. But whether two responses are considered homogeneous, and whether such homogenization is problematic, both depend on the task category. For instance, in objective math tasks, we often expect no variation in the final answer but anticipate variation in the problem-solving strategy. Whereas, for creative writing tasks, we may expect variation in key narrative components (e.g. plot, genre, setting, etc), beyond the vocabulary or embedding diversity produced by temperature-sampling. Previous work addressing output homogenization often fails to conceptualize diversity in a task-dependent way. We address this gap in the literature directly by making the following contributions. (1) We present a task taxonomy comprised of eight task categories that each have distinct concepts of output homogenization. (2) We introduce task-anchored functional diversity to better evaluate output homogenization. (3) We propose a task-anchored sampling technique that increases functional diversity for task categories where homogenization is undesired, while preserving it where it is desired. (4) We challenge the perceived existence of a diversity-quality trade-off by increasing functional diversity while maintaining response quality. Overall, we demonstrate how task dependence improves the evaluation and mitigation of output homogenization.
@@ 1 Introduction
Existing approaches to reducing output homogenization rarely take task dependence into account. Several recent works propose methods that promote diversity in the alignment process or when sampling outputs at inference-time. However, these studies often fail to conceptualize diversity in a task-specific way. For example, some methods aim to increase token-level entropy or embedding-space variation in alignment (chung2025modifying; lanchantin2025diverse; slocumdiverse; li2025jointly), while others promote diversity of viewpoints and perspectives when sampling multiple outputs (wang2025multilingual; zhang2025cultivating; zhang2025noveltybench). Without a task-dependent approach, such methods may (1) fail to encourage diversity that is meaningful for a task, and/or (2) undesirably reduce homogenization in tasks where it is desired. We address this gap in the literature directly.
We introduce a task-anchored framework to evaluate and mitigate output homogenization. We build on the notion of functional diversity (zhang2025noveltybench; shypula2025evaluating), which asks whether a user would perceive two responses as meaningfully different for a given task. We argue that LLMs should be able to conceptualize functional diversity based on the task category. Consider the stakes: if a model wrongly conceptualizes functional diversity for a task that mimics an encyclopedia inquiry, the model could misrepresent historical events in an attempt to naively reduce homogenization. Conversely, if a model wrongly conceptualizes functional diversity for a creative writing task, the model might repeat the same story arc no matter how many times a user asks the model to tell a story. We argue that task dependence should be incorporated into the way we address homogenization. Our contributions are as follows.
@@
    We introduce task-anchored functional diversity to better evaluate output homogenization (§ 3.2). In our experiments, we compare to more general diversity metrics which are not task dependent (vocabulary and embedding differences). The results show that these general metrics fail to capture task-dependent diversity. Our task-anchored metric offers an alternative evaluation approach for future studies of output homogenization.
@@
    We propose a task-anchored sampling technique to increase functional diversity (§ 3.3), improving on previous sampling methods to promote diversity (zhang2025cultivating; zhang2025noveltybench). Figure 1 offers a high-level illustration of our approach. We leverage our taxonomy to instruct models with task-dependent notions of diversity. Our approach increases functional diversity for task categories where homogenization is undesired, while preserving homogenization where it is desired (Figure 2).
@@
    We challenge the perceived existence of a diversity-quality trade-off (a common narrative in the literature) by adopting a quality measure (linwildbench; wei2025rocketeval) that also accounts for task-specific factors in quality. Figure 3 shows that the diversity-quality tradeoff prevalent in previous studies may simply be the result of mis-conceptualizing both diversity and quality. Our evaluation framework corrects both.
@@ 2.1 Homogenization in Aligned Models
Several works show that aligned LLMs exhibit output homogenization across a variety of tasks, such as creative writing (moon2024homogenizing; wu2025generative), political discussions (durmus2023towards; santurkar2023whose), and math problem-solving (slocumdiverse). zhang2024forcing; zhang2025noveltybench further show that in question-answering, models often produce the same answer, even when the question is underspecified and multiple valid answers exist. These studies often evaluate homogenization in specific task domains, suggesting that problematic notions of homogenization are task-dependent. Our proposed taxonomy compares a variety of these task categories.
Representation Concerns
@@
In some tasks, homogeneous outputs may raise concerns about representation. The literature on pluralistic alignment (sorensen2024position; chen2024pal; zhang2025cultivating) highlights representational harms, particularly when users seek advice or opinions from LLMs. However, pluralistic alignment discussions tend to operate in contexts where representation or diversity is presumed to be desirable. These discussions should recognize the task dependent nature of diversity, as we discuss in this work.
Causes of Homogenization
@@
There are many causes of output homogenization, such as limited diversity in training data and model design choices (zhang2025cultivating; fazelpour2025value). In particular, the alignment process is well-known to amplify homogenization in LLM outputs (kirk2024understanding; lanchantin2025diverse). Models are typically aligned using methods such as Reinforcement Learning with Human Feedback (RLHF) (ziegler2019fine) or Direct Preference Optimization (DPO) (rafailov2023direct). These methods involve training on a dataset of pairwise preferences $\{(x,y^{+},y^{-})\}$, where $x$ is a prompt, $y^{+}$ is a preferred response, and $y^{-}$ is a dispreferred response. When there are conflicting preferences in the training data, such that both $y^{+}\succ y^{-}\,|\,x$ and $y^{-}\succ y^{+}\,|\,x$ coexist, the RLHF and DPO objectives implicitly reward putting all sequence-level probability on the majority preference (slocumdiverse; yao2025no). Preference pairs with larger semantic differences also exert a stronger influence on the behavior of the aligned model (chung2025modifying; shen2024towards). While this line of research is important, the present work focuses on how output homogenization should be conceptualized, not why it occurs.
Outcome Homogenization
@@
Another type of homogenization occurs when multiple models produce similar outputs (kim2025correlated; wenger2025we). For example, outcome homogenization in decision-making refers to when individuals receive similar decisions from separate AI models (bommasani2022outcome; jain2024algorithmic). In this work, we focus on the single-model case and do not deal directly with homogenization across different models. But reducing homogenization within a single model likely affects homogenization across models (jain2024position; jain2025allocation).
@@ 2.2 Diversity-Promoting Methods
Alignment Methods
A growing body of literature explores methods to reduce homogenization in aligned LLMs. Several studies propose modifying the alignment process, either by altering the construction of preference datasets or by adjusting alignment objectives (lanchantin2025diverse; slocumdiverse; chung2025modifying). All of these methods substantially increase diversity during alignment, as measured by token-level entropy or embedding-space variation. However, we highlight how these metrics may not capture meaningful, task-dependent notions of diversity.
Inference-Time Methods
While most evaluations of homogenization examine temperature-sampled outputs, a few studies explore prompt-based strategies to explicitly sample diverse outputs at inference-time. For example, zhang2025cultivating use a system prompt that explicitly tells LLMs to generate $k$ responses in a single output that represent “diverse values.” zhang2025noveltybench propose in-context regeneration, where models are prompted to produce a different response while retaining all previous responses in the conversation context. Other works also use implicit techniques, such as persona-based or multilingual prompting (wang2025large; wang2025multilingual). Our work improves inference-time methods by explicitly clarifying the notion of “diversity” in model instructions.
@@ 3 Framework
Task Category Task Definition Example Task Functional
Diversity
Reward
Type
Example Dataset
A. Well-Specified
@@
Tasks with a single verifiable correct answer What is the largest
Spanish-speaking country?
None Verifiable SimpleQA
wei2024simpleQA
B. Underspecified
@@
Tasks with many verifiable correct answers Name one Spanish-
speaking country.
Different correct
@@
Verifiable (Multiple) NoveltyBench
zhang2025noveltybench
C. Random
@@
Tasks that involve randomizing over a set of options Roll a make-believe
6-sided dice.
Different pseudo-
@@
Verifiable (Multiple) NoveltyBench
zhang2025noveltybench
D. Problem-Solving
@@
Tasks to solve a problem with a verifiable solution How many positive whole-number divisors of 196? Different solution strategies Verifiable MATH-500
lightman2023let
E. Problem-Solving or Design
Subjective
Tasks to solve a problem with many partially verifiable
solutions
Design a room that minimizes energy consumption while maintaining comfort. Different solutions or strategies Partially
@@ E. Problem-Solving or Design
MacGyver
tian2024macgyver
F. Encyclopedia
@@
Tasks to provide information about real-world societies, traditions, events where there are credible references Why is Isaac Newton
famous?
Different factual
@@
Partially
Verifiable
Community Alignment
zhang2025cultivating
G. Creative Writing Tasks that require creative
@@
Non-Verifiable WildBench
linwildbench
H. Advice or
Opinions
Tasks that solicit advice,
opinions or feedback on
specific topics/scenarios
What is a good Mother’s day gift? Different views or perspectives Non-Verifiable Community Alignment
zhang2025cultivating
@@ 3.1 Task Taxonomy
We begin by outlining the task categories used in our task-anchored framework (Table 1). Each of the $8$ categories are distinguished by their conceptualization of output homogenization. The first four categories (A, B, C, D) capture prompts that elicit verifiable solution(s) that might be considered objective in nature, yet may still have more than one verifiable answer or explanation [^1] We view objective/subjective and underspecified/well-specified not as discrete categories, but as spectra that task categories span. While the terms “objective” and “subjective” invite philosophical debate, such discussions are beyond this paper’s scope. Likewise, we use “underspecified” and “well-specified” loosely, as clarifying terms, not precise definitions of the answer spaces.. The second four categories (E, F, G, H) capture prompts that elicit more open-ended solution(s) that may be considered non-verifiable or only have partially verifiable components. Our taxonomy offers a more granular categorization of reward verifiability than the binary distinction (verifiable vs non-verifiable) in the literature (lambert2024tulu; lanchantin2025bridging). Our task categories capture many real-world LLM usecases identified in recent work (tamkin2024clio; chatterji2025people) (c.f. Appendix 6.1). The categories are also motivated by recent studies that evaluate homogenization in specific task domains (§ 2.1). Though our taxonomy is non-exhaustive, we illustrate how task categories are a useful mechanism to appropriately conceptualize output homogenization.
@@
Each task category further corresponds to a specific type of response variation or functional diversity. Previous works define two responses to be functionally diverse if a user would perceive them to be meaningfully different (zhang2025noveltybench; shypula2025evaluating). In this work, we define functional diversity based on our task categories, clarifying how responses could be meaningfully different for each task. For example, in Problem Solving Objective (category D), functional diversity is in solution strategies, not in the final answer. Whereas, in Creative Writing (category G), functional diversity is in the key creative elements (e.g. plot, genre, setting, etc.), not just in character names or vocabulary.
@@ 3.2 Evaluating Task-Anchored Diversity
Next, we formalize task-anchored functional diversity. We let $\mathcal{P}$ denote the set of possible input prompts or tasks and we let $\mathcal{Y}$ denote the set of possible outputs (e.g. sequences of tokens). We adopt the simple notation that a language model $\mathcal{M}$ is a stochastic function $\mathcal{M}:\mathcal{P}\to\mathcal{Y}$ that maps each prompt $p\in\mathcal{P}$ to an output $y\in\mathcal{Y}$. For a given prompt $p$, we assume $d(p,y_{a},y_{b})\in[0,1]$ to be a pairwise diversity metric that indicates whether two responses $y_{a}$ and $y_{b}$ differ. To specify functional diversity, we anchor the definition of $d(p,y_{a},y_{b})$ on the task category $c(p)\in\mathcal{T}$, where each prompt $p$ is is associated with a task category based on Table 1.
Definition 3.1 (Task-Anchored Functional Diversity).
@@
For example, consider whether two responses are functionally diverse in the Problem-Solving Objective task (category D). Here, $d(p,y_{a},y_{b})$ represents whether responses have different solution strategies, and assumes the single verifiable answer to be the same. In practice, evaluating functional diversity requires human annotation or LLM-judges. Based on the pairwise indicators of functional diversity, a set of responses can be partitioned into functionally distinct response groups as follows.
Definition 3.2 (Number of Functionally Unique Responses).
Let $\mathcal{Y}_{p}=\{y_{1},\dots,y_{n}\}$ be a set of responses for prompt $p$. Define an undirected graph $G=(\mathcal{Y}_{p},E)$ where the edge set is $E:=\{(y_{a},y_{b})\in\mathcal{Y}_{p}\times\mathcal{Y}_{p}:d(p,y_{a},y_{b})=0\}$. Then, the number of functionally unique responses is $|\pi_{0}(G)|$ where $\pi_{0}(G)$ denotes the set of path-connected components of $G$.
Previous studies often evaluate response diversity with general diversity metrics. By general, we mean that the metric does not reference or depend on a predefined task category. For example, vocabulary diversity quantifies the extent to which two responses use different words where higher values mean more unique words or less words shared. Embedding diversity measures the difference in semantic content according to cosine distance in an embedding vector space where higher values mean more semantic difference. Appendix 6.5 provides formal definitions for these metrics.
@@
In-Context Regeneration Can you generate a
@@
Table 2: Prompt-Based Sampling Strategies. We modify prompt-based sampling methods in previous works (zhang2025cultivating; zhang2025noveltybench) to promote task-anchored functional diversity.
3.3 Promoting Task-Anchored Diversity
We introduce a task-anchored sampling technique which modifies existing prompt-based methods for promoting diversity (c.f. Figure 1). Prompt-based sampling strategies are inference-time methods to generate multiple responses. We focus on two existing methods: system prompt sampling, which generates multiple responses in a single generation (zhang2025cultivating), and in-context regeneration, which iteratively generates multiple responses (zhang2025noveltybench). Both these methods instruct the model to generate “different” or “diverse” responses. We modify these methods by clarifying in the instruction what is meant by “different” or “diverse”, using the functional diversity concepts in our taxonomy (Table 2, Appendix 6.4). To reduce homogenization at inference-time, the model could sample over these responses, or choose a response based on other alignment criteria.